OpenStudy (anonymous):

How do i set this equation up? A brownie recipe asks for one and one third times as much sugar as chocolate chips. If four and one half cups of sugar is used, what quantity of chocolate chips would then be needed, according to the recipe?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (nincompoop):

\[4\frac{ 1 }{ 2 }c \times 1\frac{ 1 }{ 3 }=b\]

OpenStudy (nincompoop):

start by converting the mixed fraction into improper fraction then perform the operation being asked (multiply)

OpenStudy (nincompoop):

I hope @ganeshie8 agrees

OpenStudy (anonymous):

so 9/2c*4/3=b ?

OpenStudy (nincompoop):

ya

OpenStudy (anonymous):

6C=B

OpenStudy (nincompoop):

that translate to 6 cups of sugar will be used for brownies

OpenStudy (anonymous):

i got it wrong 6 was correct but it needed to b in fraction form

OpenStudy (nincompoop):

were you given choices to pick from?

OpenStudy (anonymous):

no its on this website called Aleks.com

OpenStudy (nincompoop):

link me b in fraction form would just be b/1

OpenStudy (nincompoop):

no luck the link tells me it's closed

OpenStudy (anonymous):

so is there a certain formula i use to set this up A brownie recipe requires one and one third cups of sugar to one cup of chocolate chips. If three and one third cups of sugar is used, what quantity of chocolate chips will be needed, according to the recipe?

OpenStudy (nincompoop):

ya the one i showed 4/3(9/2 c) = b b is the same as b/1

OpenStudy (nincompoop):

that's the most simplification of the algebraic notation required

OpenStudy (nincompoop):

what is the answer that the website suggesting?